OpenStudy (anonymous):

Lagrange Multiplier stuff! Find the max and min values f(x,y,z)= x^2-y-z, x^2-y^2+z=0

OpenStudy (anonymous):

Gradient of f= <2x,-1,-1> Gradient of g (constraint)= <2x, -2y, 1> Now im stuck

OpenStudy (anonymous):

I did \[2x=\lambda(2x) \] \[-1=\lambda (-2y)\] \[-1=\lambda \] which doesnt make any sense cuz I have different values for lambda

OpenStudy (anonymous):

What are you assuming your second value for lambda is?

OpenStudy (anonymous):

1 because 2x=lambda (2x) So i have -1 and 1 for lambda...

OpenStudy (anonymous):

Well, my best guess is that lambda is -1 only. If we look at that 2x = lambda(2x) part: \[2x = \lambda (2x) \implies 2x-\lambda (2x) = 0 \implies 2x( 1 - \lambda) = 0\] So lambda = 1 OR x = 0. Clearly lambda must be -1, so this forces x to be 0 for this to be valid. From there, since we know lambda is -1, we can get y to be -1/2 and from there ue the constraint to see what x must be.

OpenStudy (anonymous):

To see what z must be, my bad

OpenStudy (anonymous):

oh ok. so then lambda=-1, y= 1/2, x=0, then z=1/4 f(0,1/2,1/4)=-3/4 Right? Is this the min or max?

OpenStudy (anonymous):

Right. You can use the second derivative test to see if its a min or a max.

OpenStudy (anonymous):

AWESOME:) thank you!!

OpenStudy (anonymous):

No problem :)